Emma, Jackson, and Nicholas have been selling raffle tickets to raise money for the math club. Emma sold 5 less than twice as many tickets as Nicholas. Jackson sold more than twice as many tickets as Nicholas. Together the three sold 180 tickets. How many did each student sell?
Answer by nyc_function(2741)

Here is the equation:
2x - 5 + 2x + x = 180

5x - 5 = 180

5x = 180 + 5

5x = 185

x = 185/5

x = 37

Emma sold 2(37) - 5 tickets or 69 tickets.

Jackson sold 37(2) = 74 tickets.

Nicholas sold 37 tickets.

How do we know if this is the right answer?

Let's add 37 + 74 + 69. If the total is 180, then the answer is correct.

37 + 74 + 69 = 180